A printer manufacturing company claims that it's new ink-efficient printer can print an average of 1500 pages of word documents with standard deviation of 60. Thirty-five of these printers showed a mean of 1475 pages. Does this support the company's claim? Use 95% confidence level.

1
Expert's answer
2022-06-12T15:58:45-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1500"

"H_1:\mu\not=1500"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{1475-1500}{60/\sqrt{35}}=-2.465"

Since it is observed that "|z|=2.465>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z<-2.465)=0.013701," and since "p=0.013701<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 1500, at the "\alpha = 0.05" significance level.